OpenStudy (anonymous):

Factor completely. 4x^4-4x+2xy-2y

OpenStudy (anonymous):

pls reconfirm the question

OpenStudy (anonymous):

2(2x^2 - 2x+xy-y) 2[2x(x-1)+y(x-1)] = 2(2x+y)(x-1)]

OpenStudy (anonymous):

jimmy how can 2*2x^2 be = 4x^4

OpenStudy (anonymous):

oh - right - i cant see properly today! - i took it to be 4x^2